\begin{thebibliography}{999}


%\cite{Obied:2018sgi}
\bibitem{Obied:2018sgi}
  G.~Obied, H.~Ooguri, L.~Spodyneiko and C.~Vafa,
  %``De Sitter Space and the Swampland,''
  arXiv:1806.08362 [hep-th].
  %%CITATION = ARXIV:1806.08362;%%
  %132 citations counted in INSPIRE as of 28 Dec 2018
  
    %\cite{Hertzberg:2007wc}
\bibitem{Hertzberg:2007wc}
  M.~P.~Hertzberg, S.~Kachru, W.~Taylor and M.~Tegmark,
  %``Inflationary Constraints on Type IIA String Theory,''
  JHEP {\bf 0712} (2007) 095
  doi:10.1088/1126-6708/2007/12/095
  [arXiv:0711.2512 [hep-th]].
  %%CITATION = doi:10.1088/1126-6708/2007/12/095;%%
  %150 citations counted in INSPIRE as of 28 Dec 2018
  
   %\cite{Dvali:2013eja}
\bibitem{Dvali:2013eja}
  G.~Dvali and C.~Gomez,
  %``Quantum Compositeness of Gravity: Black Holes, AdS and Inflation,''
  JCAP {\bf 1401} (2014) 023
  doi:10.1088/1475-7516/2014/01/023
  [arXiv:1312.4795 [hep-th]].
  %%CITATION = doi:10.1088/1475-7516/2014/01/023;%%
  %86 citations counted in INSPIRE as of 28 Dec 2018 
  
  
    %\cite{Dvali:2018jhn}
\bibitem{Dvali:2018jhn}
  G.~Dvali, C.~Gomez and S.~Zell,
  %``Quantum Breaking Bound on de Sitter and Swampland,''
  arXiv:1810.11002 [hep-th].
  %%CITATION = ARXIV:1810.11002;%%
  %11 citations counted in INSPIRE as of 28 Dec 2018
  
  %\cite{Agrawal:2018own}
\bibitem{Agrawal:2018own}
  P.~Agrawal, G.~Obied, P.~J.~Steinhardt and C.~Vafa,
  %``On the Cosmological Implications of the String Swampland,''
  Phys.\ Lett.\ B {\bf 784} (2018) 271
  doi:10.1016/j.physletb.2018.07.040
  [arXiv:1806.09718 [hep-th]].
  %%CITATION = doi:10.1016/j.physletb.2018.07.040;%%
  %92 citations counted in INSPIRE as of 28 Dec 2018
  
  %\cite{Denef:2018etk}
\bibitem{Denef:2018etk}
  F.~Denef, A.~Hebecker and T.~Wrase,
  %``de Sitter swampland conjecture and the Higgs potential,''
  Phys.\ Rev.\ D {\bf 98} (2018) no.8,  086004
  doi:10.1103/PhysRevD.98.086004
  [arXiv:1807.06581 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevD.98.086004;%%
  %51 citations counted in INSPIRE as of 28 Dec 2018
  
  %\cite{Murayama:2018lie}
\bibitem{Murayama:2018lie}
  H.~Murayama, M.~Yamazaki and T.~T.~Yanagida,
  %``Do We Live in the Swampland?,''
  JHEP {\bf 1812} (2018) 032
  doi:10.1007/JHEP12(2018)032
  [arXiv:1809.00478 [hep-th]].
  %%CITATION = doi:10.1007/JHEP12(2018)032;%%
  %33 citations counted in INSPIRE as of 28 Dec 2018
  
  %\cite{Choi:2018rze}
\bibitem{Choi:2018rze}
  K.~Choi, D.~Chway and C.~S.~Shin,
  %``The dS swampland conjecture with the electroweak symmetry and QCD chiral symmetry breaking,''
  JHEP {\bf 1811} (2018) 142
  doi:10.1007/JHEP11(2018)142
  [arXiv:1809.01475 [hep-th]].
  %%CITATION = doi:10.1007/JHEP11(2018)142;%%
  %25 citations counted in INSPIRE as of 28 Dec 2018
  
  %\cite{Hamaguchi:2018vtv}
\bibitem{Hamaguchi:2018vtv}
  K.~Hamaguchi, M.~Ibe and T.~Moroi,
  %``The swampland conjecture and the Higgs expectation value,''
  JHEP {\bf 1812} (2018) 023
  doi:10.1007/JHEP12(2018)023
  [arXiv:1810.02095 [hep-th]].
  %%CITATION = doi:10.1007/JHEP12(2018)023;%%
  %20 citations counted in INSPIRE as of 28 Dec 2018
 
   %\cite{Ooguri:2018wrx}
\bibitem{Ooguri:2018wrx}
  H.~Ooguri, E.~Palti, G.~Shiu and C.~Vafa,
  %``Distance and de Sitter Conjectures on the Swampland,''
  Phys.\ Lett.\ B {\bf 788} (2019) 180
  doi:10.1016/j.physletb.2018.11.018
  [arXiv:1810.05506 [hep-th]].
  %%CITATION = doi:10.1016/j.physletb.2018.11.018;%%
  %47 citations counted in INSPIRE as of 28 Dec 2018 
  
  
  %\cite{Garg:2018reu}
\bibitem{Garg:2018reu}
  S.~K.~Garg and C.~Krishnan,
  %``Bounds on Slow Roll and the de Sitter Swampland,''
  arXiv:1807.05193 [hep-th].
  %%CITATION = ARXIV:1807.05193;%%
  %57 citations counted in INSPIRE as of 28 Dec 2018
  
 
 %\cite{Symanzik:1969ek}
\bibitem{Symanzik:1969ek}
  K.~Symanzik,
  %``Renormalizable models with simple symmetry breaking. 1. Symmetry breaking by a source term,''
  Commun.\ Math.\ Phys.\  {\bf 16} (1970) 48.
  doi:10.1007/BF01645494
  %%CITATION = doi:10.1007/BF01645494;%%
  %299 citations counted in INSPIRE as of 29 Dec 2018   

%\cite{Iliopoulos:1974ur}
\bibitem{Iliopoulos:1974ur}
  J.~Iliopoulos, C.~Itzykson and A.~Martin,
  %``Functional Methods and Perturbation Theory,''
  Rev.\ Mod.\ Phys.\  {\bf 47} (1975) 165.
  doi:10.1103/RevModPhys.47.165
  %%CITATION = doi:10.1103/RevModPhys.47.165;%%
  %357 citations counted in INSPIRE as of 27 Dec 2018


%\cite{Weinberg:1987vp}
\bibitem{Weinberg:1987vp}
  E.~J.~Weinberg and A.~q.~Wu,
  %``Understanding Complex Perturbative Effective Potentials,''
  Phys.\ Rev.\ D {\bf 36} (1987) 2474.
  doi:10.1103/PhysRevD.36.2474
  %%CITATION = doi:10.1103/PhysRevD.36.2474;%%
  %198 citations counted in INSPIRE as of 12 Jan 2019

%\cite{Ringwald:1989dz}
\bibitem{Ringwald:1989dz}
  A.~Ringwald and C.~Wetterich,
  %``Average Action for the $N$ Component $\phi^4$ Theory,''
  Nucl.\ Phys.\ B {\bf 334} (1990) 506.
  doi:10.1016/0550-3213(90)90489-Z
  %%CITATION = doi:10.1016/0550-3213(90)90489-Z;%%
  %104 citations counted in INSPIRE as of 01 Jan 2019
  
  %\cite{Ooguri:2006in}
\bibitem{Ooguri:2006in}
  H.~Ooguri and C.~Vafa,
  %``On the Geometry of the String Landscape and the Swampland,''
  Nucl.\ Phys.\ B {\bf 766} (2007) 21
  doi:10.1016/j.nuclphysb.2006.10.033
  [hep-th/0605264].
  %%CITATION = doi:10.1016/j.nuclphysb.2006.10.033;%%
  %220 citations counted in INSPIRE as of 01 Jan 2019


\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{graphicx}
\usepackage{amssymb}
\usepackage{amscd}
\usepackage{amsmath}
\usepackage{cite}
\usepackage[british]{babel}
\usepackage{microtype}

\begin{document}

{\hbox to\hsize{\hfill January 2019 }}

\bigskip \vspace{3\baselineskip}

\begin{center}
{\bf \Large 
A brief remark on convexity of effective potentials and de Sitter Swampland conjectures }

\bigskip

\bigskip

{\bf Archil Kobakhidze \\ }

\smallskip

{ \small \it
School of Physics, The University of Sydney, NSW 2006, Australia, \\
E-mails: archil.kobakhidze@sydney.edu.au 
\\}

\bigskip
 
\bigskip

\bigskip

{\large \bf Abstract}

\end{center}
\noindent 
Recently proposed de Sitter Swampland conjectures imply non-trivial constraints  on  a scalar field potential in any effective field theory that admits a quantum gravity completion. The original conjecture apparently excludes many phenomenologically motivated scalar potentials with de Sitter extrema, such as the perturbative Standard Model Higgs potential and the QCD axion potential, as the viable low-energy theories. Subsequently, the refined, weaker conjecture was proposed.  However,  
the full effective potential, having been defined as a Legendre transform,  is necessarily convex. This ensures that the original Swampland conjecture can actually be satisfied in phenomenologically relevant models, providing no de Sitter vacua exist.      

\bigskip
 
\bigskip

\bigskip

Motivated by string theory considerations, it has been conjectured in  Ref. \cite{Obied:2018sgi} that that the potential $V(\phi_i)$ for scalar fields, $\phi_i$, in any low-energy effective theory that admits embedding into a fundamental theory with quantum gravity included, must satisfy the following universal bound:
\begin{equation}
\vert \partial_{\phi_i}V\vert \geq \frac{c}{M_p}\cdot V~, 
\label{1}
\end{equation}
for some positive constant $c>0$ of order 1. Here, $\vert \partial_{\phi_i}V\vert=\sqrt{\sum_i \left (\partial_{\phi_i}V\right)^2}$ computed in the basis of scalar fields with canonical kinetic terms and minimal coupling to gravity, and $M_p\approx 2.4\cdot 10^{18}$ GeV is the reduced Planck mass. Notably, the condition (\ref{1}) excludes de Sitter extrema for which $\partial_{\phi_i}V=0,~V>0$, and in particular de Sitter ground state as the origin of the observed accelerated expansion of the universe at present epoch. Hence the name, the de Sitter Swampland conjecture. Note, that the bound (\ref{1}) with $ c =\sqrt{54/13}$  has been obtained in \cite{Hertzberg:2007wc} in the context of type IIA string theory. Also, incompatibility of de Sitter vacua with quantum corpuscular description has been argued previously in \cite{Dvali:2013eja}. The relation between the corpuscular description and the conjecture (\ref{1}) has been illuminated in \cite{Dvali:2018jhn}. 

The original de Sitter Swampland conjecture of Eq. (\ref{1}) appeared to be rather restrictive for cosmological \cite{Agrawal:2018own} and particle physics \cite{Denef:2018etk, Murayama:2018lie, Choi:2018rze, Hamaguchi:2018vtv} applications. Subsequently, the weaker conjecture was proposed in \cite{Ooguri:2018wrx} (see also \cite{Garg:2018reu}), 
\begin{equation}
\vert \partial_{\phi_i}V\vert \geq \frac{c}{M_p}\cdot V~~\rm{or}~~\rm{min}\left( \partial_{\phi_i}\partial_{\phi_j}V\right) \leq -\frac{c'}{M^2_p}\cdot V~, 
\label{2}
\end{equation}  
where $c'$ is another positive constant of order 1 and  $\rm{min}\left( \partial_{\phi_i}\partial_{\phi_j}V\right)$ is a minimum eigenvalue of the Hessian matrix $\partial_{\phi_i}\partial_{\phi_j}V$. While the de Sitter minima are still prohibited by the refined conjecture, de Sitter maxima are now allowed, providing $\rm{min}\left( \partial_{\phi_i}\partial_{\phi_j}V\right)$ is sufficiently negative. 

However, if the potential $V$ is understood as the full effective potential, which is defined via a Legendre transform, it must be everywhere convex \cite{Symanzik:1969ek, Iliopoulos:1974ur}. The Hessian $\partial_{\phi_i}\partial_{\phi_j}V$ then is necessarily positive semi-definite  and hence the second inequality in Eq. (\ref{2}) becomes redundant. In other words, the strong de Sitter conjecture (\ref{1}) is automatically satisfied, providing the scalar potential admits no de Sitter minima. 

The standard perturbative 1PI effective potential fails to reflect the convexity property in a region between minima, because the perturbative calculations are based on the expansion around one definite vacuum configuration. Such constant field configuration, however, corresponds  to a unstable saddle point rather than a minimum in the spinodal region of the potential. Instead, one must take into account large-wavelength spinodal modes, an inhomogeneous mixture of minima. As a result, the effective potential $V_k$ computed at energy $k$ gets flattened in the regions between the minima as $k\to 0$ \cite{Weinberg:1987vp, Ringwald:1989dz}.  Hence,  large-scale nonperturbative fluctuations dominate near the region of the effective potential between the minima and are essential part of the consistent description of the low-energy effective theory. It is an interesting question whether these fluctuations are related to the tower of large-wavelength  states, which are discarded in \cite{Ooguri:2006in, Ooguri:2018wrx} from the spectrum of the effective theory. This may also have its bearing to the so-called field distance conjecture \cite{Ooguri:2006in}. 

To conclude, the often ignored large scale nonperturbative fluctuations of scalar fields flatten regions between minima of the effective potential and restore it convexity. Thus, these fluctuations are the essential part of the consistent low-energy description of the theory.   Therefore, the strong version of the de Sitter Swampland conjecture is automatically satisfied in phenomenologically relevant models, providing there are no de Sitter minima.

\paragraph{Acknowledgement} I acknowledge useful discussions on the Swampland conjectures with Mark Hertzberg and Tsutomu Yanagida. The work was partially supported by the Australian Research Council and the Rustaveli National Science Foundation.

\end{document}